# Mask Image Watermarking

### v1

@@ Abstract
We present MaskMark, a simple, efficient and flexible framework for image watermarking. MaskMark has two variants: MaskMark-D, which supports global watermark embedding, watermark localization, and local watermark extraction for applications such as tamper detection, and MaskMark-ED, which focuses on local watermark embedding and extraction with enhanced robustness in small regions, enabling localized image protection. Built upon the classical Encoder-Distortion-Decoder training paradigm, MaskMark-D introduces a simple masking mechanism during the decoding stage to support both global and local watermark extraction. A mask is applied to the watermarked image before extraction, allowing the decoder to focus on selected regions and learn local extraction. A localization module is also integrated into the decoder to identify watermark regions during inference, reducing interference from irrelevant content and improving accuracy. MaskMark-ED extends this design by incorporating the mask into the encoding stage as well, guiding the encoder to embed the watermark in designated local regions for enhanced robustness. Comprehensive experiments show that MaskMark achieves state-of-the-art performance in global watermark extraction, local watermark extraction, watermark localization, and multi-watermark embedding. It outperforms all existing baselines, including the recent leading model WAM for local watermarking, while preserving high visual quality of the watermarked images. MaskMark is also flexible, by adjusting the distortion layer, it can adapt to different robustness requirements with just a few steps of fine-tuning. Moreover, our approach is efficient and easy to optimize, requiring only 20 hours on a single A6000 GPU with just 1/15 the computational cost of WAM.
@@ 1 Introduction
Image watermarking (Potdar et al., 2005) is a crucial technique for embedding imperceptible information into images, serving purposes such as copyright protection, content authentication, and provenance tracking. With the proliferation of AI-generated content (AIGC) (Rombach et al., 2022; Saharia et al., 2022), the necessity of robust watermarking schemes has become even more pronounced, as distinguishing between real and synthetic images is increasingly challenging.
@@
Traditional deep image watermarking methods (Zhu et al., 2018; Tancik et al., 2020; Jia et al., 2021) typically perform global watermark embedding and extraction, treating the entire image as a uniform entity. However, this global approach suffers from several critical limitations. First, when the image undergoes heavy tampering, such as inpainting (Yu et al., 2023; Zhang et al., 2023), resulting in only a small region retaining the watermark, global methods often fail to extract it effectively. Second, even if a watermark is successfully extracted from the entire image, such methods cannot localize which region actually contains the watermark, making fine-grained forensic analysis and fair judgment difficult. Third, in scenarios where only specific regions of an image are valuable and need protection, or when different parts of the image originate from different sources and require distinct watermarking, global embedding is inherently incapable of providing the flexibility and granularity.
@@
To address these limitations, we propose MaskMark, a simple, efficient, and flexible image watermarking framework that introduces a masking mechanism during training to guide the embedding and extraction of localized watermark signals. Depending on which stage the mask is introduced, MaskMark has two variants. MaskMark-D introduces the mask only during the decoding phase. This design enables global watermark embedding while supporting localized extraction. Specifically, by applying masks over the watermarked images before watermark extraction, the decoder learns to extract watermark signals from specific regions. Meanwhile, the end-to-end training encourages the encoder to embed the watermark uniformly across the image, ensuring that all regions contain sufficient signal for reliable local extraction. Additionally, a watermark localization module is incorporated into the decoder, enabling it to identify watermarked regions during inference, which also reduces interference from irrelevant content and enhances practical robustness. MaskMark-ED introduces the mask during both the encoding and decoding phases, allowing for both localized watermark embedding and extraction. In this setting, the encoder is trained to embed not only the watermark bits but also the spatial mask into the image. This enables the model to adaptively allocate watermark strength to designated regions while leaving the rest of the image nearly untouched.
@@
Extensive experiments demonstrate that MaskMark significantly outperforms existing baselines in both global and local watermark extraction, as well as in watermark localization, while preserving image quality. Under the 32-bit embedding setting, MaskMark achieves near-perfect accuracy in global watermark extraction, even under common distortions. Notably, geometric distortions cause most baseline methods to fail completely, while MaskMark remains highly robust. For local watermark extraction, MaskMark consistently delivers strong results: when only 5% of the image carries watermark signals and no distortion is applied, it achieves nearly 100% extraction accuracy, and remains robust across a wide range of distortion scenarios. In terms of watermark localization, MaskMark demonstrates high precision, accurately identifying watermark regions even in the presence of various distortions. Furthermore, although not explicitly trained for multi-watermark embedding, MaskMark maintains high extraction and localization performance even when the number of embedded watermarks increases to five. In addition to its effectiveness, MaskMark is highly efficient. Training requires only a single A6000 GPU for approximately 20 hours, 15× less compute than the recent state-of-the-art local watermarking model WAM (Sander et al., 2025). It also scales effortlessly to higher bit lengths (e.g., 64 or 128), retaining strong local extraction performance even under distortion, whereas WAM is inherently limited to 32-bit embedding and does not scale beyond that. Furthermore, the framework supports fast fine-tuning for different applications: for example, after just 20k training steps, MaskMark can reach over 99% extraction accuracy against VAE-based adaptive attacks. These advantages make MaskMark a practical, efficient, and scalable solution for real-world watermarking applications. In summary, our main contributions are as follows:
  • We introduce MaskMark, a simple yet powerful image watermarking framework that supports both global and local embedding and extraction, along with watermark localization via a flexible masking mechanism.
  • Extensive experiments demonstrate MaskMark excels in global and local extraction, as well as multi-watermark embedding, while maintaining high visual quality of the watermarked images.
  • MaskMark is highly efficient, requiring 15× less compute than WAM, seamlessly scales to higher watermark bit lengths (e.g., 64 or 128), and supports fast fine-tuning for quick adaptation to different attack scenarios, enhancing its practicality.
@@ 2.1 Image Watermarking
Image watermarking techniques can generally be categorized into two types: global watermarking and local watermarking methods. Global watermarking methods aim to extract watermark information from the entire image. Most traditional deep learning-based approaches fall into this category. These methods focus on achieving robustness against various types of distortions, ensuring that the embedded watermark can still be reliably recovered even when the image undergoes degradation. For example, MBRS (Jia et al., 2021) specifically targets robustness against JPEG compression. Methods like StegaStamp (Tancik et al., 2020) and PIMoG (Fang & et al., 2022) are designed to handle real-world physical distortions such as screen-shooting and print-shooting. More recent approaches like ZoDiac (Zhang et al., 2024a) and SuperMark (Hu et al., 2024a) tackle adaptive attacks, while Robust-Wide (Hu et al., 2024b) and VINE (Lu et al., 2024) focus on robustness against instruction-driven image editing.
In contrast, local watermarking methods focus on extracting watermark information from a specific region of the image. Recent methods, such as WAM (Sander et al., 2025) and our proposed MaskMark, belong to this category. WAM treats watermark extraction as a segmentation task (Kirillov et al., 2023), predicting watermark bits at the pixel level. This fine-grained approach enables precise local extraction but also introduces challenges. As the predicted bit length per pixel increases (e.g., beyond 32 bits), the task becomes more difficult, causing training instability and reduced performance. Additionally, WAM’s training is computationally expensive, requiring eight V100 GPUs for nearly a week, making it impractical for widespread use. Moreover, WAM does not support native local watermark embedding; instead, it globally embeds a watermark and then crops the image to focus on the local region, introducing inherent losses during embedding that weaken the robustness of final extraction.
@@ 2.2 Watermark Localization
Watermark localization (Zhang et al., 2024b; Hu et al., 2025; Sander et al., 2025) refers to the ability to determine which regions of a watermarked image still contain watermark information after modifications. This capability enables the identification of unaltered content, serving as an active detection mechanism for tamper localization. Currently, image watermarking techniques primarily adopt two paradigms for watermark localization. The first paradigm embeds a one-dimensional copyright watermark alongside a two-dimensional localization watermark in the original image. During extraction, localization is based on the fragility of the localization watermark, which cannot be fully recovered from a modified image. Key methods in this category include EditGuard (Zhang et al., 2024b) and OmniGuard (Zhang et al., 2024c). EditGuard embeds a solid-color template within the host image and attempts to recover it from a modified version. The difference between the recovered and original templates is calculated at each pixel, and a threshold-based decision identifies watermark-preserved regions. OmniGuard improves upon EditGuard by embedding a natural image as the template, enhancing fidelity. It also introduces a Degradation-aware Tamper Extractor, improving robustness in detecting tampered regions under distortion. This paradigm requires parallel extraction of both copyright and localization watermarks, which may affect image quality. Moreover, both watermarks need independent robustness, and the presence of the template watermark does not guarantee the presence of the copyright watermark.
The second paradigm, in contrast, embeds only a one-dimensional copyright bitstream and directly determines the presence or absence of watermark information at each pixel to achieve localization. Methods such as WAM (Sander et al., 2025) and our proposed MaskMark fall under this category. WAM employs a decoder that simultaneously performs pixel-wise watermark presence detection and copyright bit extraction. In contrast, our MaskMark incorporates a dedicated localization module within the decoder, focusing solely on watermark presence detection at each pixel. This approach is more lightweight and easier to optimize. Compared to the first paradigm, this method ensures that the localized watermark regions strictly correspond to the areas containing copyright watermark information, enhancing interpretability. Additionally, it guarantees both the robustness of copyright watermark extraction and the robustness of localization.
@@ 3.1 Design Principles
In general, our objective is to enable the watermark model to effectively extract watermarks from images where only local regions are watermarked. We identify three primary reasons why traditional watermark models fail in this scenario. First, the decoder is trained exclusively on globally watermarked images and has never encountered cases where only a local region contains a watermark. Consequently, it is unable to perform zero-shot extraction when presented with such images. Second, the non-watermarked portions of an image interfere with the decoder’s extraction process, especially when the watermark occupies a small area, leading to extraction failure. Third, because the decoder is optimized for global watermark extraction, the encoder tends to dilute the watermark’s intensity over the entire image. This results in local regions having either insufficient or fragmented watermark strength, thereby exacerbating the extraction challenge.
To address these challenges, we find that simply incorporating a basic mask mechanism during the decoding stage to manipulate the watermarked image generated by the encoder is sufficient. To solve the first issue, we retain only the watermark in the regions selected by the mask while setting the pixel values of the remaining regions to zero, thereby training the decoder to extract watermarks from partially watermarked images. To overcome the second issue, we replace the non-masked regions with the original clean image and incorporate a watermark localization module in the decoder. This module learns to differentiate between watermarked and non-watermarked regions, effectively mitigating the interference from irrelevant image content. These two strategies directly address the first and second challenges and indirectly resolve the third challenge. By ensuring that the decoder can extract the watermark from local regions and accurately distinguish it from non-watermarked areas, the encoder is encouraged to distribute the watermark strength uniformly across all local regions rather than diluting it over the entire image, thereby preserving both extractability and distinctiveness.
The strategies above form the basis of our MaskMark-D. To further enhance the encoder’s ability to address the third challenge directly, we propose MaskMark-ED, which also introduces the mask mechanism during the encoding stage. In MaskMark-ED, the mask is embedded into the image along with the watermark bits during training. This enables the encoder to learn to actively concentrate the watermark within the selected regions based on the embedded mask, thereby further improving the robustness of local watermarking.
3.2 Overview
@@
Figure 1: The overall end-to-end training pipeline of MaskMark. (1) In the Mask Generation stage, we generate candidate masks from four predefined types and randomly select one mask $M$ for the subsequent stages. (2) In the Watermark Embedding stage, the encoder $\mathcal{E}$ embeds the watermark bits $W_{gt}$ into the original image $I_{orig}$, optionally using the mask $M$ (for MaskMark-ED), to produce the watermarked image $I_{wm}$. (3) In the Watermark Masking stage, the mask $M$ is used to fuse $I_{orig}$ and $I_{wm}$, resulting in the fused image $I_{fuse}$, which is then subjected to a randomly selected distortion from a predefined distortion pool, yielding the distorted image $I^{{}^{\prime}}_{fuse}$. The masked region is then cropped using $M$ to obtain $I_{mask}$. (4) In the Watermark Extraction stage, the decoder $\mathcal{D}$ extracts the predicted watermark bits $W_{pd}$ from $I_{mask}$ and the predicted mask $M_{pd}$ from $I^{{}^{\prime}}_{fuse}$.
3.3 Mask Generation
@@
The mask generation process constructs a pool of candidate masks with diverse types, from which one is randomly selected as the mask $M$ for each image during training. We follow a similar mask generation strategy as LaMa (Suvorov et al., 2022), utilizing four distinct types of masks: Full Mask, Rectangle Mask, Irregular Mask, and Segment Mask. These masks are designed to enhance the model’s ability to handle watermark embedding and extraction under diverse conditions, and each serves a specific purpose: The Full Mask enables global watermark learning across the entire image, serving as a fundamental capability. The Rectangle Mask focuses on regularly shaped local regions, encouraging the model to operate within confined areas of varying sizes. The Irregular Mask introduces complex, arbitrarily shaped regions to improve robustness in non-uniform contexts. Lastly, the Segment Mask targets semantically meaningful areas by leveraging object masks from the MS-COCO dataset (Lin et al., 2014), helping the model generalize to real-world scenarios.
3.4 Watermark Embedding
@@
We describe our method for embedding both the watermark bits $W_{gt}$ and the optional mask $M$ into the original image $I_{orig}$. The process begins by randomly sampling binary watermark bits $W_{gt}\in\{0,1\}$ of length $l$. These bits are then transformed into a feature map $f$ with shape $(C_{f},H,W)$ using a lightweight CNN, where $C_{f}$ represents the number of channels in the feature map. This CNN consists of a linear layer followed by several Conv-Norm-ReLU blocks. Specifically, the linear layer first maps $W_{gt}$ to a tensor of shape $(1,l,l)$, which is then bilinearly interpolated to match the spatial dimensions of the input image, resulting in a tensor of shape $(1,H,W)$. This tensor is then processed through the CNN to produce the feature map $f$.
To embed the watermark, we begin by concatenating the original image $I_{orig}$ with the watermark feature $f$ along the channel dimension. For the MaskMark-D variant, no further concatenation is performed, resulting in a tensor of shape $(3+C_{f},H,W)$ that encourages global watermark embedding. In contrast, the MaskMark-ED variant additionally concatenates the mask $M$, yielding a tensor of shape $(3+C_{f}+1,H,W)$. The inclusion of $M$ serves to guide the model toward localized embedding by indicating regions of interest during training. The use cases for both variants are discussed in detail in Sec. 3.8.
@@
The concatenated tensor is then passed through a U-Net (Kirillov et al., 2023) to generate an intermediate encoded image $I_{enc}$. To obtain the final watermarked image $I_{wm}$, we apply a Just-Noticeable-Difference (JND) module (Wu et al., 2017), which modulates the embedding signal based on human visual sensitivity to enhance perceptual quality:
@@
where $\mu$ is the JND scaling factor to control the watermark strength. We explore several strategies to improve the visual quality of the watermarked image and find that JND modulation consistently delivers the best performance (see Sec. 4.5 for details).
3.5 Watermark Masking
We describe how the mask $M$ is used to process the watermarked image $I_{wm}$ for subsequent mask prediction and watermark extraction by the decoder.
First, we generate a fused image $I_{fuse}$ by combining $I_{wm}$ and the original image $I_{orig}$, where the unmasked regions are replaced with the corresponding pixels from $I_{orig}$:
@@ 3.5 Watermark Masking
Next, we apply a randomly selected distortion from a predefined distortion pool to $I_{fuse}$, producing an augmented image $I^{\prime}_{fuse}$. This step follows a common practice in traditional watermarking methods to improve robustness against various transformations.
Finally, we use the mask $M$ once more to isolate the watermarked regions of $I^{\prime}_{fuse}$, setting all other pixels to zero to obtain the input $I_{mask}$ for watermark extraction:
@@
3.6 Watermark Extraction
We describe how the decoder $\mathcal{D}$ extracts the predicted mask from $I^{\prime}_{fuse}$ and recovers the watermark bits from $I_{mask}$. To achieve these two objectives, $\mathcal{D}$ consists of two dedicated modules: a U2-Net (Qin et al., 2020) for mask prediction, and a U-Net (Ronneberger et al., 2015) followed by a CNN for watermark extraction.
Specifically, the U2-Net takes $I^{\prime}_{fuse}$ as input and predicts a mask $M_{pd}$ of shape $(C_{M},H,W)$. Meanwhile, the U-Net processes $I_{mask}$ to produce an intermediate feature $f^{\prime}$ with shape $(C_{f},H,W)$, which is then passed through a CNN to obtain the predicted watermark bits $W_{pd}$. Unlike the CNN used in the encoder $\mathcal{E}$, the CNN in the decoder $\mathcal{D}$ first applies several Conv–Norm–ReLU layers to further extract features from $f^{\prime}$. The resulting features are then interpolated to a fixed shape of $(1,l,l)$ and subsequently transformed by linear layers into a bit sequence of length $l$, yielding the final watermark prediction $W_{pd}$.
3.7 Training Objectives
@@
$\mathcal{L}_{\text{enc}}=\mathcal{L}_{\text{MSE}}(I_{\text{wm}},I_{\text{orig}}).$ (4)
Note that we impose constraints only in the pixel space, as we found that this setup, combined with JND modulation, already achieves high visual quality. Introducing constraints in the feature space or using GAN-based losses would negatively impact the overall performance, as discussed in Sec. 4.5.
The decoder loss is formulated as:
$\mathcal{L}_{\text{dec}}=$ $\ \mathcal{L}_{\text{MSE}}(W_{\text{pd}},W_{\text{gt}})+\alpha\mathcal{L}_{\text{MSE}}(M_{\text{pd}},M),$ (5)
@@
$\mathcal{L}_{\text{total}}=\beta_{enc}\cdot\mathcal{L}_{\text{enc}}+\beta_{dec}\cdot\mathcal{L}_{\text{dec}},$ (6)
where $\beta_{enc}$ and $\beta_{dec}$ are the weights for the encoder and decoder losses, respectively.
Compared to conventional watermarking methods, our approach introduces only a mask loss at the decoder stage. As a result, the method retains a simple yet effective objective, making it easy to extend.

### v2

@@ Abstract
We present MaskMark, a simple, efficient, and flexible framework for image watermarking. MaskMark has two variants: (1) MaskMark-D, which supports global watermark embedding, watermark localization, and local watermark extraction for applications such as tamper detection; (2) MaskMark-ED, which focuses on local watermark embedding and extraction, offering enhanced robustness in small regions to support fine-grined image protection. MaskMark-D builds on the classical encoder-distortion layer-decoder training paradigm. In MaskMark-D, we introduce a simple masking mechanism during the decoding stage that enables both global and local watermark extraction. During training, the decoder is guided by various types of masks applied to watermarked images before extraction, helping it learn to localize watermarks and extract them from the corresponding local areas. MaskMark-ED extends this design by incorporating the mask into the encoding stage as well, guiding the encoder to embed the watermark in designated local regions, which improves robustness under regional attacks. Extensive experiments show that MaskMark achieves state-of-the-art performance in global and local watermark extraction, watermark localization, and multi-watermark embedding. It outperforms all existing baselines, including the recent leading model WAM for local watermarking, while preserving high visual quality of the watermarked images. In addition, MaskMark is highly efficient and adaptable. It requires only 20 hours of training on a single A6000 GPU, achieving 15× computational efficiency compared to WAM. By simply adjusting the distortion layer, MaskMark can be quickly fine-tuned to meet varying robustness requirements.
@@ 1 Introduction
Image watermarking [14] is a crucial technique for embedding imperceptible information into images, serving purposes such as copyright protection, content authentication, and provenance tracking. With the proliferation of AI-generated content (AIGC) [16; 18], the boundary between real and synthetic images has become increasingly blurred, making it especially important to develop robust watermarking schemes for content verification and traceability.
@@
Traditional deep image watermarking methods [30; 21; 9] typically perform global watermark embedding and extraction, treating the entire image as a uniform entity. However, these global approaches suffer from several critical limitations. 1. Lack of local watermark extraction: When an image undergoes heavy tampering, such as inpainting [24; 26], the watermark may survive only in a small, local region that remains untouched by manipulation. In such cases, global methods often fail to extract the watermark effectively. 2. Inability to localize the watermark: Even if a watermark is successfully extracted from the image, global methods cannot localize which region actually contains the watermark, making it difficult for fine-grained forensic analysis and fair judgment. 3. Lack of local watermark embedding: In scenarios where only specific regions of an image are valuable and need protection, or when different parts of the image originate from different sources and require distinct watermarking, global embedding is inherently incapable of providing the flexibility and granularity.
We argue that the training paradigm of traditional global watermarking methods, which treats the entire watermarked image as a whole for both encoding and decoding, prevents the encoder and decoder from developing spatial awareness. Specifically, the decoder cannot identify which regions of the image contain watermark signals for effective extraction, while the encoder lacks the ability to adaptively embed the watermark into specific spatial regions.
@@
Based on the above analysis, we propose MaskMark, a simple, efficient, and flexible image watermarking framework. MaskMark introduces a masking mechanism during training to guide the model learn spatially aware embedding and extraction of local watermark signals. Depending on the stage which the mask is introduced, MaskMark offers two variants: MaskMark-D introduces the mask only during the decoding phase, enabling global watermark embedding while supporting local extraction. Specifically, by applying masks to retain only the selected regions of the watermarked images during extraction, the decoder is guided to identify which regions contain watermark signals and to focus on them for effective local extraction. MaskMark-ED introduces the mask during the encoding and decoding phases, allowing the embedding and extraction of the local watermark. In this setting, the encoder is trained to embed not only the watermark bits but also the spatial mask into the image. This allows the encoder to leverage the mask to adaptively allocate watermark strength to designated regions, while keeping the rest of the image nearly unaffected.
@@
Extensive experiments demonstrate that MaskMark significantly outperforms existing baselines in both global and local watermark extraction, watermark localization, and local watermark embedding, while preserving image quality. Specifically, for local watermark extraction, MaskMark achieves a nearly 100% extraction accuracy even when only 5% of the image carries watermark signals. In terms of watermark localization, MaskMark demonstrates high precision in identifying watermark regions. Furthermore, although not explicitly trained for multi-watermark embedding, MaskMark maintains strong extraction and localization performance even when embedding up to 5 distinct watermarks in a single image. More importantly, MaskMark exhibits strong robustness across a wide range of distortions, including geometric distortions that typically break many existing watermarking methods. In addition to its effectiveness, MaskMark is highly efficient. Training the encoder-decoder model requires only approximately 20 hours on a single A6000 GPU, which is 15× less compute than the recent state-of-the-art local watermarking model WAM [19]. MaskMark also scales effortlessly to different bit lengths (e.g., 32, 64, and 128), whereas WAM is inherently limited to 32-bit embedding and does not scale beyond that. Furthermore, MaskMark supports fast fine-tuning for different use cases. For example, MaskMark can reach a near 100% extraction accuracy against VAE-based adaptive attacks after just 20k training steps. These advantages make MaskMark a practical, efficient, and scalable solution for real-world applications.
@@ 2.1 Image Watermarking
Image watermarking techniques can generally be categorized into two types: global watermarking and local watermarking methods. Global watermarking methods aim to extract watermark information from the entire image. Most traditional deep learning-based approaches fall into this category. These methods focus on achieving robustness against various types of distortions, ensuring that the embedded watermark can still be reliably recovered even when the image undergoes degradation. For example, MBRS [9] specifically targets robustness against JPEG compression. Methods like StegaStamp [21] and PIMoG [4] are designed to handle real-world physical distortions such as screen-shooting and print-shooting. More recent approaches like ZoDiac [25] and SuperMark [5] tackle adaptive attacks, while Robust-Wide [7] and VINE [13] focus on robustness against instruction-driven image editing.
In contrast, local watermarking methods focus on extracting watermark information from a specific region of the image. Recent methods, such as WAM [19] and our proposed MaskMark, belong to this category. WAM treats watermark extraction as a segmentation task [11], predicting watermark bits at the pixel level and then averaging these per-pixel predictions to obtain the final result. While this fine-grained approach enables local watermark extraction, it also presents several challenges. 1. Limited extraction from small regions: When the watermarked area is very small, only a few pixels contribute to extraction, making naive averaging unreliable, especially under distortions. 2. Lack of scalability with longer messages: WAM struggles to scale beyond 32-bit messages, as training becomes increasingly difficult with longer bit lengths. 3. High computational cost: Training WAM is resource-intensive, requiring eight V100 GPUs for nearly a week, which limits its practicality. 4. Lack of native local embedding: WAM embeds watermarks globally and then crops for local focus, introducing embedding losses that reduce extraction robustness.
@@ 2.2 Watermark Localization
Watermark localization [28; 6; 19] refers to the ability to determine which regions of a watermarked image still contain watermark information after modifications. This capability enables the identification of unaltered content, serving as an active detection mechanism for tamper localization. Currently, image watermarking techniques primarily adopt two paradigms for watermark localization. The first paradigm embeds a one-dimensional copyright watermark alongside a two-dimensional localization watermark in the original image. During extraction, localization is based on the fragility of the localization watermark, which cannot be fully recovered from a modified image. Key methods in this category include EditGuard [28] and OmniGuard [29]. EditGuard embeds a solid-color template within the host image and attempts to recover it from a modified version. The difference between the recovered and original templates is calculated at each pixel, and a threshold-based decision identifies watermark-preserved regions. OmniGuard improves upon EditGuard by embedding a natural image as the template, enhancing fidelity. It also introduces a Degradation-aware Tamper Extractor, improving robustness in detecting tampered regions under distortion. This paradigm requires parallel extraction of both copyright and localization watermarks, which may affect image quality. Moreover, both watermarks need independent robustness, and the presence of the template watermark does not guarantee the presence of the copyright watermark.
The second paradigm, in contrast, embeds only a one-dimensional copyright bitstream and directly determines the presence or absence of watermark information at each pixel to achieve localization. Methods such as WAM [19] and our proposed MaskMark fall under this category. WAM employs a decoder that simultaneously performs pixel-wise watermark presence detection and copyright bit extraction. In contrast, our MaskMark incorporates a dedicated localization module within the decoder, focusing solely on watermark presence detection at each pixel. This approach is more lightweight and easier to optimize. Compared to the first watermark localization paradigm, this method ensures that the local watermarked regions strictly correspond to the areas containing copyright watermark information, enhancing interpretability. Additionally, it guarantees both the robustness of copyright watermark extraction and the robustness of localization.
@@ 3.1 Design Principles
In general, we have three main objectives: local watermark extraction, watermark localization, and local watermark embedding. Among these, our primary goal is local watermark extraction, which aims to recover the embedded message from images where only a small, spatially local region contains the watermark signal. In practice, we find that achieving high performance on this task naturally necessitates solving the other two problems as well. We now identify three key reasons why traditional watermark models fail under this setting.
First, since the decoder is trained exclusively on globally watermarked images and has never encountered cases with only locally embedded watermarks, it fails to perform zero-shot extraction on such inputs. Second, the non-watermarked portions of an image interfere with the decoder’s extraction process, especially when the watermark occupies a small area, leading to extraction failure. Third, because the decoder is optimized for global watermark extraction, the encoder tends to dilute the watermark’s intensity over the entire image. This results in local regions having either insufficient or fragmented watermark strength, thereby exacerbating the extraction challenge.
To address these challenges, we propose MaskMark-D, which introduces a basic mask mechanism during the decoding stage to guide the decoder in identifying and focusing on watermark-containing regions. To solve the first issue, we retain the watermark only in the masked regions and set other regions’ pixels to zero, training the decoder to extract watermarks from partially watermarked images. For the second issue, we replace the non-masked regions with the original clean image and add a watermark localization module in the decoder to differentiate between watermarked and non-watermarked areas, reducing interference from irrelevant content. These two strategies directly address the first two issues by enabling the decoder to extract watermarks from local regions, while also indirectly mitigating the third issue by guiding the encoder to evenly distribute the watermark under the end-to-end training, thus facilitating the decoder’s extraction process.
To further enhance the encoder’s ability to address the third challenge, we propose MaskMark-ED, which incorporates the mask during the encoding stage to explicitly guide watermark placement. In MaskMark-ED, the mask is embedded into the image along with the watermark bits during training. This enables the encoder to learn to actively concentrate the watermark within the selected regions based on the embedded mask, thereby further improving the robustness of local watermarking.
3.2 Training
@@
Figure 1: The overall end-to-end training pipeline of MaskMark. (1) In the Mask Generation stage, we generate candidate masks from four predefined types and randomly select one mask $M$ for the subsequent stages. (2) In the Watermark Embedding stage, the encoder $\mathcal{E}$ embeds the watermark bits $W_{gt}$ into the original image $I_{orig}$, optionally using the mask $M$ (for MaskMark-ED), to produce the watermarked image $I_{wm}$. (3) In the Watermark Masking stage, the mask $M$ is used to fuse $I_{orig}$ and $I_{wm}$ (see Eq. 2), resulting in the fused image $I_{fuse}$, which is then subjected to a randomly selected distortion from a predefined distortion pool, yielding the distorted image $I^{{}^{\prime}}_{fuse}$. The masked region is then cropped using $M$ to obtain $I_{mask}$. (4) In the Watermark Extraction stage, the decoder $\mathcal{D}$ extracts the predicted mask $M_{pd}$ from $I^{{}^{\prime}}_{fuse}$ and the predicted watermark bits $W_{pd}$ from $I_{mask}$.
Mask Generation.
@@
The mask generation process constructs a pool of candidate masks with diverse types, from which one mask $M$ is randomly selected for each image during training. We follow a similar mask generation strategy as LaMa [20] for image inpainting, utilizing four distinct types of masks: Full Mask, Rectangle Mask, Irregular Mask, and Segment Mask. These masks can enhance the model’s ability to handle watermark embedding and extraction under diverse conditions from different perspectives: Full Mask enables the model to embed and extract watermarks across the entire image, serving as a fundamental capability. Rectangle Mask focuses on regularly shaped local regions, encouraging the model to operate within confined areas of varying sizes. Irregular Mask introduces complex, arbitrarily shaped regions to improve robustness in non-uniform contexts. Lastly, Segment Mask targets semantically meaningful areas by leveraging object masks from the MS-COCO dataset [12], helping the model generalize to real-world scenarios.
Watermark Embedding.
@@
We describe the process of embedding both the watermark bits $W_{gt}$ and the optional mask $M$ into the original image $I_{orig}$. We first randomly sample binary watermark bits $W_{gt}\in{0,1}$ of length $l$, which are transformed into a feature map $f\in\mathbb{R}^{C_{f}\times H\times W}$ using a lightweight CNN, where $C_{f}$ is the number of channels in $f$, and $H$ and $W$ denote the height and width of $I_{orig}$, respectively. This CNN consists of a linear layer that maps $W_{gt}$ to a tensor of shape $(1,l,l)$, followed by bilinear interpolation to $(1,H,W)$ and several Conv-Norm-ReLU (CNR) blocks that produce the final feature map $f$.
To embed the watermark, we concatenate the original image $I_{orig}$ with the watermark feature $f$ along the channel dimension. For the MaskMark-D, this results in a tensor of shape $(3+C_{f},H,W)$, promoting global watermark embedding. For the MaskMark-ED, the optional mask $M$ is further concatenated, yielding a tensor of shape $(3+C_{f}+1,H,W)$. The mask guides the model to focus on the selected regions during training, enabling local embedding within those specified areas.
@@
The concatenated tensor is then passed through a U-Net [11] to generate an intermediate encoded image $I_{enc}$. To obtain the final watermarked image $I_{wm}$, we apply a Just-Noticeable-Difference (JND) module [22], which modulates the embedding signal based on human visual sensitivity to enhance the perceptual quality:
@@
where $\mu$ is the JND scaling factor to control the watermark strength. We explore several strategies to improve the visual quality of the watermarked image and find that JND modulation consistently delivers the best performance (see Appendix D.1 for details).
Watermark Masking.
We describe how the mask $M$ is used to process the watermarked image $I_{wm}$ for subsequent mask prediction and watermark extraction by the decoder. First, we generate a fused image $I_{fuse}$ by combining $I_{wm}$ and the original image $I_{orig}$, where the unmasked regions are replaced with the corresponding pixels from $I_{orig}$:
@@
Next, a randomly selected distortion from a predefined distortion pool is applied to $I_{fuse}$, producing an augmented image $I^{\prime}_{fuse}$. This step follows a common practice in traditional watermarking methods to improve robustness against various transformations. Finally, we use the mask $M$ once more to isolate the watermarked regions of $I^{\prime}_{fuse}$, setting all other pixels to zero to obtain the input $I_{mask}$ for watermark extraction:
@@
Watermark Extraction.
We describe how the decoder $\mathcal{D}$ extracts the predicted mask from $I^{\prime}_{fuse}$ and recovers the watermark bits from $I_{mask}$. To achieve these two objectives, $\mathcal{D}$ consists of two dedicated modules: a U2-Net [15] for mask prediction, and a U-Net [17] followed by a CNN for watermark extraction. Specifically, the U2-Net takes $I^{\prime}_{fuse}$ as input and predicts a mask $M_{pd}$. Meanwhile, the U-Net processes $I_{mask}$ to produce an intermediate feature $f^{\prime}$ with shape $(C_{f},H,W)$, which is then passed through a CNN to obtain the predicted watermark bits $W_{pd}$. Unlike the CNN used in the encoder $\mathcal{E}$, the CNN in the decoder $\mathcal{D}$ first applies several Conv–Norm–ReLU layers to further extract features from $f^{\prime}$. The resulting features are then interpolated to a fixed shape of $(1,l,l)$ and subsequently transformed by linear layers into a bit sequence of length $l$, yielding the final watermark prediction $W_{pd}$.
Training Objectives.
@@
$\mathcal{L}_{\text{enc}}=\mathcal{L}_{\text{MSE}}(I_{wm},I_{orig}).$ (4)
Note that we impose constraints only in the pixel space, as we find that this setup, combined with JND modulation, already achieves high visual quality. Introducing constraints in the feature space or using GAN-based losses would negatively impact the overall performance, as discussed in Appendix D.1. The decoder loss is formulated as:
$\mathcal{L}_{\text{dec}}=$ $\ \mathcal{L}_{\text{MSE}}(W_{pd},W_{gt})+\alpha\mathcal{L}_{\text{MSE}}(M_{pd},M),$ (5)
@@
$\mathcal{L}_{\text{total}}=\beta_{\text{enc}}\cdot\mathcal{L}_{\text{enc}}+\beta_{\text{dec}}\cdot\mathcal{L}_{\text{dec}},$ (6)
where $\beta_{\text{enc}}$ and $\beta_{\text{dec}}$ are the weights for the encoder and decoder losses, respectively. Compared to conventional watermarking methods, our approach introduces only a mask loss at the decoder stage. As a result, it retains a simple yet effective objective, making it easy to extend.
3.3 Inference
Figure 2: Watermark localization and extraction process in our decoder during inference.
The inference process consists of three main stages: watermark embedding, localization, and extraction. Embedding is performed by the encoder. In MaskMark-D, the encoder embeds the watermark bits across the entire image. In MaskMark-ED, users provide both the watermark bits and a mask, allowing the encoder to embed the watermark into specific regions. Localization and extraction are performed by the decoder. As shown in Figure 2, unlike conventional global extraction methods, our approach first uses the decoder’s localization module to identify watermark-containing regions. The rest of the image is masked out by setting non-watermarked areas to zero, reducing noise during extraction. The watermark is then recovered from the retained regions. As discussed in Sec. 4.5, this strategy improves extraction robustness, especially when only small regions remain watermarked.